OpenStudy (anonymous):

hat is the approximate value of on gigabyte or 2 to the power of 30 bytes?

jhonyy9 (jhonyy9):

do you know these ? 8 bits = 1 byte (B) 1024 bytes = 1 Kilobyte (KB) 1024 Kilobytes = 1 Megabyte (MB) 1024 Megabytes = 1 Gigabyte (GB) 1024 Gigabytes = 1 Terabyte (TB) or 2^30 = ? 2^10 =1024 2^30 = 1024*1024*1024 = ?

OpenStudy (anonymous):

Thanks